# roundedstyle backgroundcolor=verylightgray, roundcorner=10pt, innertopmargin=10pt, innerbottommargin=10pt, innerleftmargin=10pt, innerrightmargin=10pt, linewidth=1pt

### v2

@@ 5 Results
Main results (RQ1). Table 1 and 2 present the results of self-correction with the guidelines generated by MAGIC on the development set of BIRD and SPIDER, respectively. Overall, the guidelines generated by our proposed method, MAGIC, outperform all the baselines in terms of total effectiveness measured by execution accuracy and demonstrate significant improvements in self-correction performance across the BIRD and SPIDER datasets, outperforming self-correction guidelines written by expert humans. For example, MAGIC’s guidelines improve the DIN-SQL Pourreza et al. (2023) baseline from 56.52 to 59.13 in terms of execution accuracy and outperform the self-correction guideline of Wang et al. (2024a) as well, as presented in Tables 5 and 1 respectively.
We experiment with self-correction in three scenarios, as shown in Table 1: (i) correcting incorrect SQLs, assuming the availability of a correctness oracle. This setup is common in recent studies focusing on self-correction in text-to-SQL Chen et al. (2023); Zhang et al. (2024). (ii) SQLs with execution errors: this scenario is common in self-correction where there is no available correctness oracle, but the goal is to avoid generating incorrect and non-executable SQL Wang et al. (2024a). (iii) All SQLs: this setup is common in methods where self-correction is part of a multi-step model, employing a self-correction guideline on all predictions Pourreza et al. (2023).
@@
Self-Consistency Wang et al. (2022) 81.64
Multiple-Prompt Lee et al. (2024) 80.22
@@
Human’s G Wang et al. (2024a) 81.15
Human’s G Pourreza et al. (2023) 80.35
@@
Table 2: Effectiveness results for self-correcting incorrect SQL queries on the SPIDER development set. We compared MAGIC’s generated guideline (MAGIC’s G) against two top self-correction baselines: self-consistency and multiple-prompt, as well as two expert human self-correction guidelines.
@@
We find that applying self-correction to all predicted SQLs, as in the human-expert written guideline for self-correction in Wang et al. (2024a), reduces effectiveness. Intuitively, when self-correction is applied for all predicted SQLs, there is a risk that the LLM changes its previously correct prediction to an incorrect one, as suggested by previous work Li et al. (2024d). Therefore, designing a proper self-correction guideline is crucial, or it is safer to apply self-correction where a correctness oracle can identify whether the initial response is correct or not. This oracle can be a human user who minimally interacts with the system and only determines whether the executed query results meet their expectations. Regarding the "SQLs with execution errors" scenario, only 46 queries predicted by the initial text-to-SQL system Pourreza et al. (2023) are non-executable, which limits the effectiveness of self-correction for these cases, with several methods achieving the same effectiveness (49.03) for medium query difficulty.
@@
It is noteworthy that although we have not attempted to combine MAGIC with other baselines, as our focus is on analyzing different self-correction methods rather than integrating them, some baselines could potentially enhance MAGIC if combined, such as Self-consistency Wang et al. (2022) and Multiple-Prompt Lee et al. (2024).
@@
Feedback generation effectiveness (RQ3). Figure 3 presents an analysis of the number of corrected SQL queries with and without the inclusion of the manager agent, by evaluating the performance of the MAGIC framework across a range of maximum iteration counts used as stopping criteria, from one to five. When the manager agent is excluded, the feedback agent controls the maximum iteration count and utilizes the SQL executor tool to compare the execution results of the revised SQL with the ground truth SQL. This comparison underscores the significant role of the manager agent in correcting a larger number of incorrect SQL queries within fewer iterations and ultimately correcting more SQL queries when both setups reach the maximum of five iterations. It is important to note that excluding the manager agent from MAGIC is equivalent to the critic-refine cycle method Kamoi et al. (2024), that has been employed in other tasks, such as code generation. However, unlike MAGIC, the critic-refine Kamoi et al. (2024) does not produce any guidelines as a result of its feedback generation process and has not been applied to text-to-SQL tasks to the best of our knowledge.
@@
Zero-shot GPT-4 OpenAI (2023b) 40.18
@@
Few-shot GPT-4 Li et al. (2023) 45.66
Few-shot GPT-4 Li et al. (2023) + MAGIC G (Ours) 50.38
DIN-SQL Pourreza et al. (2023) 56.52
DIN-SQL Pourreza et al. (2023) + MAGIC G (Ours) 59.13
Table 5: The effectiveness of self-correction guideline generated by our proposed method, MAGIC, across different methods.
Figure 4: Analyzing impact of MAGIC’s correction guideline across different databases of development set of BIRD dataset.
We further analyze the impact of manager agent in terms of overall performance of self-correction considering all the queries. Table 4 measures the effectiveness of MAGIC in terms of execution accuracy (EX) to determine how many could be addressed correctly in total with the already corrected predictions by LLM and self-corrected ones with MAGIC. We also measure Valid Efficiency Score (VES) to determine whether the self-correction by MAGIC results in optimized SQLs compared to the ground truth. As it can be observed, MAGIC with the manager agent outperforms the exclusion of manager agent in all setups in terms of all metrics across the BIRD and SPIDER datasets. Results show that MAGIC with manager agent is able to generate more efficient SQLs compared to exclusion of manager agent, e.g., 84.19 vs. 79.58 in terms of VES for w/ manager and w/o manager on the BIRD dataset where the maximum numebr of iteration is set to 5. We manually analyzed the revised prompt by manager agent and found out during its interactions with the agents where it revises the prompts, the manager agent tends add instructions for self-correction to optimize the efficiency of SQL, which might be the reason behind generating more efficient SQLs.
6 Discussion
Applicability to other methods. We analyze the applicability of the guideline generated by MAGIC based on the mistakes of the DIN-SQL Pourreza et al. (2023) method on two other methods that generate SQL in a single step: zero-shot GPT-4 OpenAI (2023a) and few-shot GPT-4 Li et al. (2023) as shown in Table 5. As we can observe, MAGIC’s guidelines improve the zero-shot GPT-4 baseline from 40.18 to 48.19 and few-shot GPT-4 from 45.56 to 50.38 in terms of execution accuracy, respectively.
@@ 7 Conclusions
8 Limitations
In this paper, we have demonstrated that generative LLMs are capable of generating their own self-correction guidelines even more effectively than exsiting guidelines written by human experts in previous studies. However, it is important to highlight that such training data might not be available for the new query languages. Moreover, while our experiments illustrate the superior effectiveness of our self-correction guidelines in comparison to those developed by human experts, we have not conducted a systematic study to quantify the potential generation of hallucinated data during this process and its subsequent impact on the performance of the automatically generated guidelines.
Additionally, despite the high effectiveness of the generated guidelines, we wish to underscore that for large businesses, it may still be needed to have humans review these guidelines. Nevertheless, given that the generated guidelines tend not to be overly lengthy and can be generated under 2 hours (presented in Figure 11 to Figure 16 in appendix section), we believe that the effort required for reviewing the content of these guidelines is significantly more efficient than manually analyzing all of the LLM’s errors and developing self-correction guidelines from scratch.

### v3

@@ Results
Main results (RQ1). Table 1 and 2 present the results of self-correction with the guidelines generated by MAGIC on the development set of BIRD and SPIDER, respectively. Overall, the guidelines generated by our proposed method outperform all the baselines in terms of total effectiveness measured by execution accuracy and demonstrate significant improvements in self-correction performance across the BIRD and SPIDER datasets, outperforming self-correction guidelines written by expert humans. For example, MAGIC’s guidelines improve the DIN-SQL (Pourreza and Rafiei 2023) baseline from 56.52 to 59.13 in terms of execution accuracy and outperform the self-correction guideline of (Wang et al. 2024a) as well, as presented in Tables 5 and 1 respectively.
We experiment with self-correction in three scenarios, as shown in Table 1: (i) correcting incorrect SQLs, assuming the availability of a correctness oracle. This setup is common in recent studies focusing on self-correction in text-to-SQL (Chen et al. 2023; Zhang et al. 2024). (ii) SQLs with execution errors: this scenario is common in self-correction where there is no available correctness oracle, but the goal is to avoid generating incorrect and non-executable SQL (Wang et al. 2024a). (iii) All SQLs: this setup is common in methods where self-correction is part of a multi-step model, employing a self-correction guideline on all predictions (Pourreza and Rafiei 2023).
@@
Self-Consistency (Wang et al. 2022) 81.64
Multiple-Prompt (Lee et al. 2024) 80.22
@@
Human’s G (Wang et al. 2024a) 81.15
Human’s G (Pourreza and Rafiei 2023) 80.35
@@
Table 2: Effectiveness results for self-correcting incorrect SQL queries on the SPIDER development set. We compared MAGIC’s generated guideline (MAGIC’s G) against two top self-correction baselines: self-consistency and multiple-prompt, as well as expert human self-correction guidelines.
@@
We find that applying self-correction to all predicted SQLs, as in the human-expert written guideline for self-correction in (Wang et al. 2024a), reduces effectiveness. Intuitively, when self-correction is applied for all predicted SQLs, there is a risk that the LLM changes its previously correct prediction to an incorrect one, as suggested by previous work (Li et al. 2024e). Therefore, designing a proper self-correction guideline is crucial, or it is safer to apply self-correction where a correctness oracle can identify whether the initial response is correct or not. This oracle can be a human user who minimally interacts with the system and only determines whether the executed query results meet their expectations. Regarding the "SQLs with execution errors" scenario, only 46 queries predicted by the initial text-to-SQL system (Pourreza and Rafiei 2023) are non-executable, which limits the effectiveness of self-correction for these cases, with several methods achieving the same effectiveness (49.03) for medium query difficulty. It is noteworthy that although we have not attempted to combine MAGIC with other baselines, as our focus is on analyzing different self-correction methods rather than integrating them, some baselines could potentially enhance MAGIC if combined, such as Self-consistency (Wang et al. 2022) and Multiple-Prompt (Lee et al. 2024).
@@
Feedback generation effectiveness (RQ3). Figure 7 presents an analysis of the number of corrected SQL queries with and without the inclusion of the manager agent, by evaluating the performance of the MAGIC framework across a range of maximum iteration counts used as stopping criteria, from one to five. When the manager agent is excluded, the feedback agent controls the maximum iteration count and utilizes the SQL executor tool to compare the execution results of the revised SQL with the ground truth SQL. This comparison underscores the significant role of the manager agent in correcting a larger number of incorrect SQL queries within fewer iterations and ultimately correcting more SQL queries when both setups reach the maximum of five iterations. It is important to note that excluding the manager agent from MAGIC is equivalent to the critic-refine cycle method (Kamoi et al. 2024), that has been employed in other tasks, such as code generation. However, unlike MAGIC, the critic-refine (Kamoi et al. 2024) does not produce any guidelines as a result of its feedback generation process and has not been applied to text-to-SQL tasks to the best of our knowledge.
@@
Zero-shot GPT-4 (OpenAI 2023b) 40.18
@@
Few-shot CoT GPT-4 (Li et al. 2023) 45.66
Few-shot CoT GPT-4 (Li et al. 2023) + MAGIC G (Ours) 50.38
DIN-SQL (Pourreza and Rafiei 2023) 56.52
DIN-SQL (Pourreza and Rafiei 2023) + MAGIC G (Ours) 59.13
MAC-SQL (Wang et al. 2024a) 59.39
MAC-SQL (Wang et al. 2024a) + MAGIC G (Ours) 61.92
Table 5: The effectiveness of self-correction guideline generated by MAGIC across different methods.
We further analyze the impact of manager agent in terms of overall performance of self-correction considering all the queries. Table 4 measures the effectiveness of MAGIC in terms of execution accuracy (EX) to determine how many could be addressed correctly in total with the already corrected predictions by LLM and self-corrected ones with MAGIC. We also measure Valid Efficiency Score (VES) to determine whether the self-correction by MAGIC results in optimized SQLs compared to the ground truth. As it can be observed, MAGIC with the manager agent outperforms the exclusion of manager agent in all setups in terms of all metrics across the BIRD and SPIDER datasets. Results show that MAGIC with manager agent is able to generate more efficient SQLs compared to exclusion of manager agent, e.g., 84.19 vs. 79.58 in terms of VES for w/ manager and w/o manager on the BIRD dataset where the maximum number of iteration is set to 5. We manually analyzed the revised prompt by manager agent and found out during its interactions with the agents where it revises the prompts, the manager agent tends to add instructions for self-correction to optimize the efficiency of SQL, which might be the reason behind generating more efficient SQLs.
Figure 8: Analyzing impact of MAGIC’s correction guideline across different databases of BIRD dataset.
Discussion
Applicability to other methods. We analyze the applicability of the guideline generated by MAGIC based on the mistakes of the DIN-SQL (Pourreza and Rafiei 2023) method on three other methods that generate SQL in a single step: zero-shot GPT-4 (OpenAI 2023a), few-shot CoT GPT-4 (Li et al. 2023), and multi-agent (MAC-SQL (Wang et al. 2024a)) as shown in Table 5. As we can observe, MAGIC’s guidelines improve the zero-shot GPT-4 baseline from 40.18 to 48.19, Few-shot CoT GPT-4 from 45.56 to 50.38, and multi-agent baseline from 59.39 to 61.92 in terms of execution accuracy, respectively.
@@ Conclusions
Appendix A Prompts
The prompts that are used are as follow:
  • The prompt for the predefined interaction of manager agent with feedback agent can be seen in Figure 9, and the prompt with that the manager agent revises its previous predefined prompt can be seen in Figure 11.
  • The prompt for the predefined interaction of manager agent with feedback agent can be seen in Figure 10, and the prompt with that the manager agent revises its previous predefined prompt can be seen in Figure 12.
  • The prompt for manager agent to generate the self-correction guideline can be seen in Figure 13.
Furthermore, an example of last iteration of our proposed method, MAGIC, on a question of BIRD dataset, that led to a successful self-correction can be seen in Figure 14.
Appendix B Generated Self-correction guideline by MAGIC
The self-correction guideline generated by our proposed method, MAGIC, automatically and from scratch, is presented in Figures 15, 16, 17, 18, 19, and 20.
{mdframed}
[backgroundcolor=verylightgray,roundcorner=10pt] System role prompt:
Complete the text in chat style like a database manager expert. Write in simple present without using correct SQL. Accept what the user identifies as correct or incorrect.
User role prompt:
"question": "{question}",
"evidence": "{evidence}",
"Correct SQL": "{Correct SQL}",
"Incorrect SQL": "{Incorrect SQL}",
Incorrect SQL mistakes are:
Figure 9: The predefined prompt for interaction of magic agent with feedback agent.
{mdframed}
[backgroundcolor=verylightgray,roundcorner=10pt] System role prompt:
Your task is to correct the predicted SQL based on the provided feedback by expert human.
1. Input Information: You will receive a question, a database schema, a predicted SQL query, and a human feedback.
2. SQL format in your response:
- You must ensure that your response contains a valid SQL.
- The format of SQL in your response must be in the following format: “‘sql SQL “‘.
User role prompt:
- Schema Overview: {schema}
- Columns Description: columns Description
- Question: {question}
- Predicted SQL: “‘sql {Incorrect SQL} “‘
- Expert Human Feedback: {feedback}
Figure 10: The predefined prompt for interaction of magic agent with correction agent.
{mdframed}
[backgroundcolor=verylightgray,roundcorner=10pt] System role prompt
You are a helpful AI assistant that manages other assistants.
User role prompt
Manager, please review the following prompt for the following agent and generate a revised prompt.
So far you have revised the prompt for {iteration_number} times.
Agent description: This agent generates feedback based on the comparison between predicted and ground truth SQL queries.
Previous output of agent that was not useful: {agent_outputs[-1]}
Previous prompt of agent that you should revise: {agent_prompts[-1]}
Revise prompt (Return the entire prompt so it can be directly pass to the agent):
Figure 11: The prompt template of manager for revising the predefined prompts of feedback agent.
{mdframed}
[backgroundcolor=verylightgray,roundcorner=10pt] System role prompt:
You are a helpful AI assistant that manages other assistants.
User role prompt:
Manager, please review the following prompt for the following agent and generate a revised prompt.
So far you have revised the prompt for {iteration_number} times.
Agent description: This agent generates corrections for SQL queries based on expert human feedback.
Previous output of agent that was not useful: {agent_outputs[-1]}
Previous prompt of agent that you should revise: {agent_prompts[-1]}
The ideal output of agent should be the following but we cannot directly give the ideal output to the agent:
Ideal output of agent: {Correct SQL}
Revise prompt (Return the entire prompt so it can be directly pass to the agent):
Figure 12: The prompt template of manager for revising the predefined prompts of correction agent.
{mdframed}
[backgroundcolor=verylightgray,roundcorner=10pt] # Guideline format:
number. Reminder of mistake
- Question: "Question"
- Incorrect SQL generated by me: “‘sql incorrect sql “‘
- Corrected SQL generated by me: “‘sql corrected sql “‘
- Negative and strict step-by-step ask-to-myself questions to prevent same mistake again:
# Guideline so far:
{current_guideline} // current guideline variable is empty at first
# Recent mistakes that must be aggregate to Guideline:
{batch_of_successful_feedbacks}
# Updated Guideline (Return the entire of guideline):
Figure 13: The prompt of manager for self-correction guideline generation.
Figure 14: Illustration of last iteration of our proposed method, MAGIC, on a question of BIRD dataset, that led to a successful self-correction.
Figure 15: Self-correction guideline that is generated by our method, MAGIC, automatically. Page 1.
Figure 16: Self-correction guideline that is generated by our method, MAGIC, automatically. Page 2.
Figure 17: Self-correction guideline that is generated by our method, MAGIC, automatically. Page 3.
Figure 18: Self-correction guideline that is generated by our method, MAGIC, automatically. Page 4.
Figure 19: Self-correction guideline that is generated by our method, MAGIC, automatically. Page 5.
Figure 20: Self-correction guideline that is generated by our method, MAGIC, automatically. Page 6.